an iodized salt contains 0.5% NaI. a person takes 3g of salt everyday. the number of iodide ions going into his body everyday?

Grade:11

2 Answers

Vikas TU
7 years ago
(150/x)*100 = 0.5
x = 15000/0.5 = 30000 gm
 
If 150 gm of NaI is present in 30000 gm of salt.
then 3 gm of salt will have let y gm of NaI.
thus,
from unitary method,
y = 150*3/30,000 = 0.015 gm.
No of I ions =  (0.015/127)*Na = 0.711377 * 10^20 atoms. 
Ritu
6 years ago
The answer which is provided above is wrong .its right answer is 6.022*10¹9 .Mass of NaI = 150 According to question 0.5*3/100/150 (given mass/molecular mass)= 1/10*10*10*10Moles *Na =mass1/1000*6.022*10²³ =6.022*10¹9